A passenger train leaves a train depot 2 hours after a freight train leaves the same depot. The freight train is traveling 20 mph slower than the passenger train. Find the rate of each train if the passenger train overtakes the freight train in 3 hours.
Answer by mananth(16946)

freight train = x mph
passenger train = x+20 mph
...
freight train time = 3 +2 = 5 hours
passanger train = 3 hours
..
distance same. because they meet
..
5x=3(x+20)
5x=3x+60
5x-3x=60
2x= 60
/2
x= 30mph speed of freight train
passenger train speed = 30+20 = 50mph